# Free multiderivations of connected subgraph arrangements

### v2

@@ 2.2. Connected subgraph arrangements
We summarize the important definitions and results from [CK24]. We start by recalling the definition of a connected subgraph arrangement. In [CK24] the authors introduced this concept for an arbitrary underlying field. Here we restrict ourselves to arrangements over ${\mathbb{Q}}$.
Definition 2.17 ([CK24, Def. 1.1]).
@@
[…]
  • The cycle-graph $C_{n}$ on $n$ vertices. Draw $n$ vertices and connect vertex $1$ and $2$, $2$ and $3$, $\dots$, $n$ and $1$.
[…]
  • The path-with-triangle-graph $\Delta_{n,k}$, where $11\leq k\leq n-$ 1. Draw a path-graph on $n$ vertices, add an additional vertex $n+1$, and connect the vertices $k$ and $n+1$ by an edge as well as the vertices $k+1$ and $n+1$.
[…]
In [Y09], M. Yoshinaga investigated the freeness question of multiplicities for a certain class of arrangements. His setting turns out to be important for our study as well.
Definition 2.19 ([Y09, Def. 2.1]).
@@
    $A_{i_{1}}\dot{\cup}A_{i_{2}}=A_{i_{3}}$ for some choice of pairwise distinct indices $1\leq i_{1},i_{2},i_{3}\leq 3$ where $\dot{\cup}$ denotes a disjoint set union.
@@ 2.3. Extensions of a multiarrangement
The following definitions and results are due to Yoshinaga [Y09].
@@
Let $({\mathscr{A}},\mu)$ be a multiarrangement in $\mathbb{C}^{\ell}$.
@@
    $({\mathscr{A}},\mu)$ is is extendable if it can be obtained as a Ziegler-restriction of a free arrangement in $\mathbb{C}^{\ell+1}$.
@@
One of the main results in [Y09] is the following.
Theorem 2.25 ([Y09, Thm. 2.5]).
@@ 2.4. Locally and globally mixed products
In [ATW08], a very useful tool is introduced to determine whether a given multiarrangement $({\mathscr{A}},\mu)$ fails to be free.
@@
    If the localization $({\mathscr{A}}_{X},\mu_{X})$ is free with exponents $\exp({\mathscr{A}}_{X},\mu_{X})=(d_{1}^{X},d_{2}^{X},\dots,d_{k}^{X})$ for each $X\in L_{k}$, then define the $k$-th local mixed product by
    $\operatorname{LMP}(k)=\sum_{X\in L_{k}}d_{1}^{X}d_{2}^{X}\cdots d_{k}^{X}.$
@@
    If $({\mathscr{A}},\mu)$ is free with exponents $\exp({\mathscr{A}},\mu)=(d_{1},\dots,d_{\ell})$, define the $k$-th global mixed product by
    $\operatorname{GMP}(k)=\sum d_{i_{1}}d_{i_{2}}\cdots d_{i_{k}},$
@@
Theorem 2.28 ([ATW07, Cor. 4.6]).
If $({\mathscr{A}},\mu)$ is free with $\exp({\mathscr{A}},\mu)=(d_{1},\dots,d_{\ell})$ then for every $0\leq k\leq\ell$, we have
@@
For $1\leq i\leq\ell$, let $d_{i}\in\mathbb{N}$. We use the shorthand $(d_{1},\ldots,d_{\ell})_{\leq}\in\mathbb{N}^{\ell}$ to mean $d_{1}\leq d_{2}\leq\ldots\leq d_{\ell}$. Then for the latter define the $k$-th global mixed product by
$\operatorname{GMP}(k):=\sum d_{i_{1}}d_{i_{2}}\cdots d_{i_{k}},$
where the sum is taken over all $k$-tuples from $(d_{1},\ldots,d_{\ell})_{\leq}$ with $1\leq i_{1}<\dots<i_{k}\leq\ell$.
@@
Following [ATW07], given two ordered sequences of non-negative integers $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ and $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$ as above satisfying $\sum d_{i}=\sum e_{i}$, we call $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ more balanced than $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$ if $d_{\ell}-d_{1}<e_{\ell}-e_{1}$. It follows from [ATW07, p.836] that if $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ is more balanced than $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$, then $\operatorname{GMP}(k)$ is larger for $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ than for $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$.
@@
With notation as in Definition 2.29 let $S_{1}=(4,4,4,4)_{\leq}$ and $S_{2}=(2,4,5,5)_{\leq}$. Then $S_{1}$ is more balanced than $S_{2}$ since $4-4<5-2$. Among all tuples $(d_{1},d_{2},d_{3},d_{4})_{\leq}$ with $\sum d_{i}=16$, we see that $S_{1}$ is most balanced.
The following is immediate from Theorem 2.28 and Remark 2.30.
@@
    Let $\mu$ be a multiplicity on ${\mathscr{A}}$ and let $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$ be the unique most balanced sequence of integers with sum equal to $|\mu|$. If $\operatorname{GMP}(2)$ for $(e_{1},\ldots,e_{\ell})_{\leq}$ is smaller than $\operatorname{LMP}(2)$ for $({\mathscr{A}},\mu)$, then $({\mathscr{A}},\mu)$ is not free.
@@
    If ${\mathscr{A}}$ is free, then $1\in\exp({\mathscr{A}})$. Let $(e_{2},\dots,e_{\ell})_{\leq}$ be the unique most balanced sequence of integers with sum equal to $|{\mathscr{A}}|-1$. If $\operatorname{GMP}(2)$ for $(1,e_{2},\ldots,e_{\ell})_{\leq}$ is smaller than $\operatorname{LMP}(2)$ for ${\mathscr{A}}$, then ${\mathscr{A}}$ is not free.
@@ 3. Free connected subgraph multiarrangements
Let $G$ be one of the graphs $G_{3}$ to $G_{20}$ listed in Figures 1 and 2, then the corresponding connected subgraph arrangement ${\mathscr{A}}_{G}$ is totally non-free.
@@
Let $G$ be one of the graphs $G_{3}$ to $G_{20}$ listed in Figures 1 and 2. For each such $G$ we present an element $X\in L({\mathscr{A}}_{G})$ in Table 1 such that the corresponding localization $({\mathscr{A}}_{G})_{X}$ is a generic arrangement with $\operatorname{rank}({\mathscr{A}}_{G})_{X}<|({\mathscr{A}}_{G})_{X}|$. By Theorems 2.1 and 2.6 we can conclude that ${\mathscr{A}}_{G}$ is totally non-free.
@@
$G_{3}$ $H_{1}\cap H_{123}\cap H_{145}$ $3$ $\{H_{1},H_{123},H_{145},H_{12345}\}$
$G_{4}$ $H_{1}\cap H_{124}\cap H_{135}$ $3$ $\{H_{1},H_{124},H_{135},H_{12345}\}$
$G_{5}$ $H_{1}\cap H_{124}\cap H_{135}$ $3$ $\{H_{1},H_{124},H_{135},H_{12345}\}$
$G_{6}$ $H_{12}\cap H_{14}\cap H_{1235}$ $3$ $\{H_{12},H_{14},H_{1235},H_{1345}\}$
$G_{7}$ $H_{23}\cap H_{136}\cap H_{1235}\cap H_{1245}$ $4$ $\{H_{23},H_{136},H_{1235},H_{1245},H_{12346}\}$
$G_{8}$ $H_{1}\cap H_{146}\cap H_{1245}\cap H_{12367}\cap H_{12467}$ $5$ $\{H_{1},H_{146},H_{1245},H_{12367},H_{12467},H_{123456}\}$
$G_{9}$ $H_{1}\cap H_{124}\cap H_{135}$ $3$ $\{H_{1},H_{124},H_{135},H_{12345}\}$
$G_{10}$ $H_{12}\cap H_{23}\cap H_{1245}$ $3$ $\{H_{12},H_{23},H_{1245},H_{2345}\}$
$G_{11}$ $H_{13}\cap H_{23}\cap H_{1345}$ $3$ $\{H_{13},H_{23},H_{1345},H_{2345}\}$
$G_{12}$ $H_{23}\cap H_{34}\cap H_{1235}$ $3$ $\{H_{23},H_{34},H_{1235},H_{1345}\}$
$G_{13}$ $H_{15}\cap H_{35}\cap H_{1245}$ $3$ $\{H_{15},H_{35},H_{1245},H_{2345}\}$
$G_{14}$ $H_{15}\cap H_{35}\cap H_{1245}$ $3$ $\{H_{15},H_{35},H_{1245},H_{2345}\}$
$G_{15}$ $H_{125}\cap H_{135}\cap H_{245}$ $3$ $\{H_{125},H_{135},H_{245},H_{345}\}$
$G_{16}$ $H_{15}\cap H_{35}\cap H_{1245}$ $3$ $\{H_{15},H_{35},H_{1245},H_{2345}\}$
$G_{17}$ $H_{24}\cap H_{25}\cap H_{1234}$ $3$ $\{H_{24},H_{25},H_{1234},H_{1235}\}$
$G_{18}$ $H_{13}\cap H_{14}\cap H_{1235}$ $3$ $\{H_{13},H_{14},H_{1235},H_{1245}\}$
$G_{19}$ $H_{125}\cap H_{135}\cap H_{245}$ $3$ $\{H_{125},H_{135},H_{245},H_{345}\}$
$G_{20}$ $H_{15}\cap H_{35}\cap H_{1245}$ $3$ $\{H_{15},H_{35},H_{1245},H_{2345}\}$
Table 1. Elements $X\in L({\mathscr{A}}_{G_{i}})$ such that $({\mathscr{A}}_{G_{i}})_{X}$ is generic
@@
Conversely, assume that we have a graph $G=(E,V)$ and a multiplicity $\mu\geq\mathbbm{1}$ on ${\mathscr{A}}_{G}$ such that $({\mathscr{A}}_{G},\mu)$ is free.
@@
2. Assume that $G$ has at least $2$ cycles of length at least $3$ that share an edge. Since we assume that $|V|\geq 5$, the graph cannot be equal to $G_{1}$ or $G_{2}$. Let $C_{1},C_{2}$ be cycles of $G$ with a shared edge $(v_{1},v_{2})$, choose $w_{1}\in C_{1},w_{2}\in C_{2}$ such that $w_{1}\not\in C_{2},w_{2}\not\in C_{1}$. Assume first that both cycles have length $3$. In this case the induced graph on $\{v_{1},v_{2},w_{1},w_{2}\}$ is of type $G_{1}$ or $G_{2}$. Since $|V|\geq 5$ and $G$ is connected there exists a vertex $v$ that is a neighbor of at least one element of $\{v_{1},v_{2},w_{1},w_{2}\}$, while not belonging to either $C_{1}$ or $C_{2}$. Taking the induced subgraph on $\{v_{1},v_{2},w_{1},w_{2},v\}$ yields one of $G_{9}$ up to $G_{19}$. So ${\mathscr{A}}_{G}$ is again totally non-free, thanks to Proposition 3.1. Assume that at least $C_{1}$ is of length greater than $3$. Choose $v_{1},v_{2},w_{1},w_{2}$ as in the first part above and an additional $w_{3}\in C_{1}$ and take the induced subgraph on $\{v_{1},v_{2},w_{1},w_{2},w_{3}\}$. Now contract each edge at least one vertex of which is not in $\{v_{1},v_{2},w_{1},w_{2},w_{3}\}$. We now arrive at one of $G_{11},G_{12},G_{14},G_{15},G_{16},G_{17},G_{19},G_{20}$. So ${\mathscr{A}}_{G}$ is totally non-free, again by Proposition 3.1.
@@
We note that ${\mathscr{A}}_{G_{2}}$ is the famous counterexample by Edelmann and Reiner given in [ER93] to Orlik’s conjecture. In this context ${\mathscr{A}}_{G_{2}}$ appears as a non-free restriction of a free arrangement ${\mathscr{B}}$. As a multiarrangement $({\mathscr{A}}_{G_{2}},\mu_{2})=({\mathscr{B}}^{\prime\prime},\kappa)$ is the Ziegler restriction of ${\mathscr{B}}$ (recall Theorem 2.2). This gives an alternative proof of the freeness of $({\mathscr{A}}_{G_{2}},\mu_{2})$. Note that ${\mathscr{A}}_{G_{2}}={\mathscr{A}}_{G_{1}}\cup\{H_{24}\}$. It turns out that starting with the inductively free multiarrangement $({\mathscr{A}}_{G_{1}},\mu_{1})$ and adding $H_{24}$ with multiplicity $1$ gives an inductively free multiplicity on ${\mathscr{A}}_{G_{2}}$.

### v3

@@ 2.2. Connected subgraph arrangements
We summarize the important definitions and results from [5]. We start by recalling the definition of a connected subgraph arrangement. In [5] the authors introduced this concept for an arbitrary underlying field. Here we restrict ourselves to arrangements over $\mathbb{Q}$}.
Definition 2.17 ([5, Def. 1.1]).
@@
    $A_{n,k}$:1122kknnn+1n+1
  • The cycle-graph $C_{n}$ on $n$ vertices. Draw $n$ vertices and connect vertex $1$ and $2$, 22 and $3$, $\dots$, $n$ and $1$.
    $C_{n}$:1122n1n-1nn
  • The path-with-triangle-graph $\Delta_{n,k}$, where $1\leq k\leq n-1$. Draw a path-graph on $n$ vertices, add an additional vertex $n+1$, and connect the vertices $k$ and $n+1$ by an edge as well as the vertices $k+1$ and $n+1$.
    $\Delta_{n,k}$:1122kkk+1k+1nnn+1n+1
In [18], M. Yoshinaga investigated the freeness question of multiplicities for a certain class of arrangements. His setting turns out to be important for our study as well.
Definition 2.19 ([18, Def. 2.1]).
@@
[…]
A property for arrangements is called local if it is preserved under localizations, [5, Def. 6.1]. Typical examples of local properties of arrangements are freeness [11, Thm. 4.37] and inductive freeness [10, Thm 1.1]. We record two graph theoretic constructions from [5] which allow us to control local properties of the ${\mathscr{A}}_{G}$.
@@ 2.3. Extensions of a multiarrangement
The following definitions and results are due to Yoshinaga [18].
@@
Let $({\mathscr{A}},\mu)$ be a free multiarrangement in $\mathbb{C}^{\ell}$.
@@
    $({\mathscr{A}},\mu)$ is extendable if it can be obtained as a Ziegler restriction of a free arrangement in $\mathbb{C}^{\ell+1}$.
@@
One of the main results in [18] is the following.
Theorem 2.25 ([18, Thm. 2.5]).
@@
Note that a free multiarrangement $({\mathscr{A}},\mu)$ need not be extendable in general, [18, Ex. 2.8].
@@ 2.4. Locally and globally mixed products
In [4], a very useful tool is introduced to determine whether a given multiarrangement $({\mathscr{A}},\mu)$ fails to be free.
@@
    If the localization $({\mathscr{A}}_{X},\mu_{X})$ is free with exponents $\exp({\mathscr{A}}_{X},\mu_{X})=(e_{1}^{X},e_{2}^{X},\dots,e_{k}^{X})$ for each $X\in L({\mathscr{A}})_{k}$, then define the $k$-th local mixed product by
    $\operatorname{LMP}(k)=\sum_{X\in L_{k}}e_{1}^{X}e_{2}^{X}\cdots e_{k}^{X}.$
@@
    If $({\mathscr{A}},\mu)$ is free with exponents $\exp({\mathscr{A}},\mu)=(e_{1},\dots,e_{\ell})$, define the $k$-th global mixed product by
    $\operatorname{GMP}(k)=\sum e_{i_{1}}e_{i_{2}}\cdots e_{i_{k}},$
@@
Theorem 2.28 ([3, Cor. 4.6]).
If $({\mathscr{A}},\mu)$ is free with $\exp({\mathscr{A}},\mu)=(e_{1},\dots,e_{\ell})$ then for every $0\leq k\leq\ell$, we have
@@
For $1\leq i\leq\ell$, let $e_{i}\in\mathbb{N}$. We use the shorthand $(e_{1},\ldots,e_{\ell})_{\leq}\in\mathbb{N}^{\ell}$ for $e_{1}\leq e_{2}\leq\ldots\leq e_{\ell}$. Then for the latter define the $k$-th global mixed product by
$\operatorname{GMP}(k):=\sum e_{i_{1}}e_{i_{2}}\cdots e_{i_{k}},$
where the sum is taken over all $k$-tuples from $(e_{1},\ldots,e_{\ell})_{\leq}$ with $1\leq i_{1}<\dots<i_{k}\leq\ell$.
@@
Following [3], given two ordered $\ell$-tuples of non-negative integers, $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ and $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$ as above satisfying $\sum d_{i}=\sum e_{i}$, we call $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ more balanced than $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$ if $d_{\ell}-d_{1}<e_{\ell}-e_{1}$. It follows from [3, p.836] that if $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ is more balanced than $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$, then $\operatorname{GMP}(k)$ is larger for $(d_{1},d_{2},\dots,d_{\ell})_{\leq}$ than for $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$.
@@
With notation as in Definition 2.29 let $S_{1}=(4,4,4,4)_{\leq}$ and $S_{2}=(2,4,5,5)_{\leq}$. Then $S_{1}$ is more balanced than $S_{2}$ since 44<524-4<5-2. Among all quadruples $(d_{1},d_{2},d_{3},d_{4})_{\leq}$ with $\sum d_{i}=16$, we see that $S_{1}$ is most balanced.
The following is immediate from Theorem 2.28 and Remark 2.30. With respect to part (ii) of the following lemma, recall that $1\in\exp({\mathscr{A}})$ if $\mathscr{A}$} is free.
@@
    Let $\mu$ be a multiplicity on $\mathscr{A}$} and let $(e_{1},e_{2},\dots,e_{\ell})_{\leq}$ be the unique most balanced ordered $\ell$-tuple of integers with sum equal to $|\mu|$. If the $\operatorname{GMP}(2)$ for $(e_{1},\ldots,e_{\ell})_{\leq}$ is smaller than the $\operatorname{LMP}(2)$ for $({\mathscr{A}},\mu)$, then $({\mathscr{A}},\mu)$ is not free.
@@
    Let $(e_{2},\dots,e_{\ell})_{\leq}$ be the unique most balanced ordered $\ell$-tuple of integers with sum equal to $|{\mathscr{A}}|-1$. If the $\operatorname{GMP}(2)$ for $(1,e_{2},\ldots,e_{\ell})_{\leq}$ is smaller than the $\operatorname{LMP}(2)$ for $\mathscr{A}$}, then $\mathscr{A}$} is not free.
@@ 3. Free connected subgraph multiarrangements
Let $G$ be one of the graphs $G_{3}$ to $G_{20}$ listed in Figures 1 and 2. Then the corresponding connected subgraph arrangement ${\mathscr{A}}_{G}$ is totally non-free.
@@
For each $G$ as in the statement above, we present an element $X\in L({\mathscr{A}}_{G})$ in Table 1 such that the corresponding localization $({\mathscr{A}}_{G})_{X}$ is a generic arrangement with $\operatorname{rank}({\mathscr{A}}_{G})_{X}<|({\mathscr{A}}_{G})_{X}|$. Consequently, owing to Theorems 2.1 and 2.6 we conclude that ${\mathscr{A}}_{G}$ is totally non-free. ∎
@@
$G_{3}$ $H_{1}\cap H_{123}\cap H_{145}$ 33 $\{H_{1},H_{123},H_{145},H_{12345}\}$
$G_{4}$ $H_{1}\cap H_{124}\cap H_{135}$ 33 $\{H_{1},H_{124},H_{135},H_{12345}\}$
$G_{5}$ $H_{1}\cap H_{124}\cap H_{135}$ 33 $\{H_{1},H_{124},H_{135},H_{12345}\}$
$G_{6}$ $H_{12}\cap H_{14}\cap H_{1235}$ 33 $\{H_{12},H_{14},H_{1235},H_{1345}\}$
$G_{7}$ $H_{23}\cap H_{136}\cap H_{1235}\cap H_{1245}$ 44 $\{H_{23},H_{136},H_{1235},H_{1245},H_{12346}\}$
$G_{8}$ $H_{1}\cap H_{146}\cap H_{1245}\cap H_{12367}\cap H_{12467}$ 55 $\{H_{1},H_{146},H_{1245},H_{12367},H_{12467},H_{123456}\}$
$G_{9}$ $H_{1}\cap H_{124}\cap H_{135}$ 33 $\{H_{1},H_{124},H_{135},H_{12345}\}$
$G_{10}$ $H_{12}\cap H_{23}\cap H_{1245}$ 33 $\{H_{12},H_{23},H_{1245},H_{2345}\}$
$G_{11}$ $H_{13}\cap H_{23}\cap H_{1345}$ 33 $\{H_{13},H_{23},H_{1345},H_{2345}\}$
$G_{12}$ $H_{23}\cap H_{34}\cap H_{1235}$ 33 $\{H_{23},H_{34},H_{1235},H_{1345}\}$
$G_{13}$ $H_{15}\cap H_{35}\cap H_{1245}$ 33 $\{H_{15},H_{35},H_{1245},H_{2345}\}$
$G_{14}$ $H_{15}\cap H_{35}\cap H_{1245}$ 33 $\{H_{15},H_{35},H_{1245},H_{2345}\}$
$G_{15}$ $H_{125}\cap H_{135}\cap H_{245}$ 33 $\{H_{125},H_{135},H_{245},H_{345}\}$
$G_{16}$ $H_{15}\cap H_{35}\cap H_{1245}$ 33 $\{H_{15},H_{35},H_{1245},H_{2345}\}$
$G_{17}$ $H_{24}\cap H_{25}\cap H_{1234}$ 33 $\{H_{24},H_{25},H_{1234},H_{1235}\}$
$G_{18}$ $H_{13}\cap H_{14}\cap H_{1235}$ 33 $\{H_{13},H_{14},H_{1235},H_{1245}\}$
$G_{19}$ $H_{125}\cap H_{135}\cap H_{245}$ 33 $\{H_{125},H_{135},H_{245},H_{345}\}$
$G_{20}$ $H_{15}\cap H_{35}\cap H_{1245}$ 33 $\{H_{15},H_{35},H_{1245},H_{2345}\}$
Table 1. Elements $X\in L({\mathscr{A}}_{G_{i}})$ such that $({\mathscr{A}}_{G_{i}})_{X}$ is generic.
@@
Conversely, assume that we have a graph $G=(N,E)$ and a multiplicity $\mu\geq\mathbbm{1}$ on ${\mathscr{A}}_{G}$ such that $({\mathscr{A}}_{G},\mu)$ is free.
@@
2. Assume that $G$ has at least $2$ cycles of length at least $3$ that share an edge. Since we assume that $|N|\geq 5$, the graph cannot be equal to $G_{1}$ or $G_{2}$. Let $C_{1},C_{2}$ be cycles of $G$ with a shared edge $(v_{1},v_{2})$, choose $w_{1}\in C_{1},w_{2}\in C_{2}$ such that $w_{1}\not\in C_{2},w_{2}\not\in C_{1}$. Assume first that both cycles have length $3$. In this case the induced graph on $\{v_{1},v_{2},w_{1},w_{2}\}$ is of type $G_{1}$ or $G_{2}$. Since $|N|\geq 5$ and $G$ is connected there exists a vertex $v$ that is a neighbor of at least one element of $\{v_{1},v_{2},w_{1},w_{2}\}$, while not belonging to either $C_{1}$ or $C_{2}$. Taking the induced subgraph on $\{v_{1},v_{2},w_{1},w_{2},v\}$ yields one of $G_{9}$ up to $G_{19}$. So ${\mathscr{A}}_{G}$ is again totally non-free, thanks to Proposition 3.1. Assume that at least $C_{1}$ is of length greater than $3$. Choose $v_{1},v_{2},w_{1},w_{2}$ as in the first part above and an additional $w_{3}\in C_{1}$ and take the induced subgraph on $\{v_{1},v_{2},w_{1},w_{2},w_{3}\}$. Now contract each edge at least one vertex of which is not in $\{v_{1},v_{2},w_{1},w_{2},w_{3}\}$. We now arrive at one of $G_{11},G_{12},G_{14},G_{15},G_{16},G_{17},G_{19},G_{20}$. So ${\mathscr{A}}_{G}$ is totally non-free, again by Proposition 3.1.
@@
We note that ${\mathscr{A}}_{G_{2}}$ appears in the famous counterexample by Edelmann and Reiner [8] to Orlik’s conjecture as a non-free restriction of a free arrangement $\mathscr{B}$}. As a multiarrangement $({\mathscr{A}}_{G_{2}},\mu_{2})=({\mathscr{B}}^{\prime\prime},\kappa)$ is the Ziegler restriction of $\mathscr{B}$} (recall Theorem 2.2). This gives an alternative proof of the freeness of $({\mathscr{A}}_{G_{2}},\mu_{2})$. Note that ${\mathscr{A}}_{G_{2}}={\mathscr{A}}_{G_{1}}\cup\{H_{24}\}$. It turns out that starting with the inductively free multiarrangement $({\mathscr{A}}_{G_{1}},\mu_{1})$ and adding $H_{24}$ with multiplicity $1$ gives an inductively free multiplicity on ${\mathscr{A}}_{G_{2}}$.